# Amandine Brunetto  Sascha Hornauer  Fabien Moutarde Centre for Robotics

### v1

@@ Abstract
Sound plays a major role in human perception, providing essential scene information alongside vision for understanding our environment. Despite progress in neural implicit representations, learning acoustics that match a visual scene is still challenging. We propose NeRAF, a method that jointly learns acoustic and radiance fields. NeRAF is designed as a Nerfstudio module for convenient access to realistic audio-visual generation. It synthesizes both novel views and spatialized audio at new positions, leveraging radiance field capabilities to condition the acoustic field with 3D scene information. At inference, each modality can be rendered independently and at spatially separated positions, providing greater versatility. We demonstrate the advantages of our method on the SoundSpaces dataset. NeRAF achieves substantial performance improvements over previous works while being more data-efficient. Furthermore, NeRAF enhances novel view synthesis of complex scenes trained with sparse data through cross-modal learning. Project page: https://amandinebtto.github.io/NeRAF
@@ 1 Introduction
Sound is fundamental to human perception. Picture yourself in the heart of a bustling city center. It is not just what you see, but also the cacophony of car horns, footsteps, and chatter that guide your perception and decision-making. Beyond providing information about our environment, sound offers major insights into context and atmosphere, enriching our understanding of the world in ways that sight alone cannot reach. In gaming and AR/VR simulations, sound plays a crucial role in someone’s immersion, to match the environment and make things feel real.
@@
Recent advances in novel view synthesis allowed the generation of high-quality photorealistic images at any camera position from a set of captured images [34, 33, 62, 53]. They opened many exciting applications for simulated environments. Yet these methods lack acoustic synthesis essential to immersive experiences. Learning scene acoustics presents additional challenges [30]. Indeed, sound is influenced by the geometry of the space and the materials of objects and surfaces. As sound travels from its source to our ears, it undergoes multiple acoustic phenomena. Room Impulse Responses (RIR) capture them all relative to pairs of listener and sound source positions. However, acquiring RIRs is a challenging and laborious task that involves to play and record sounds from multiple source-listener positions. Consequently, recent research [31, 47, 48] focuses on estimating RIRs at novel poses from sparsely collected data. Drawing inspiration from NeRF success in image synthesis, emerging works [30, 51, 28] learn acoustic properties of the environment using neural fields, enabling the synthesis of sound from new source and microphone locations. Still, these methods do not sufficiently account for the scene geometry and materials, despite their major role in acoustics.
In response, we introduce NeRAF, a method that synthesizes high-fidelity spatialized audios and enhances NeRF novel-view synthesis (Figure 1). NeRAF is designed as a Nerfstudio module [53] that jointly learns radiance and acoustic fields. We leverage the radiance field to obtain a 3D voxel-grid containing scene color and density. We encode it to condition our neural acoustic field with implicit scene geometry and materials. Similar to previous works [30, 51, 28], we validate the effectiveness of our method on Sound Spaces dataset [6, 9]. Our contributions can be summarized as follow:
  • NeRAF improves binaural RIR predictions on the SoundSpaces dataset by $17.4\%$ T60, $31.3\%$ EDT and $31.6\%$ C50, enabling more realistic perceptual experiences.
  • Our method is data-efficient, outperforming the current state-of-the-art with only half the audio recordings.
  • NeRAF supports spatially separate renderings of images and audios.
  • We alleviate the constraints on the training data by allowing the use of non-co-located microphones and cameras.
  • NeRAF enhances novel-view synthesis in complex scenes with sparse training data through joint learning of acoustic and radiance fields.
  • We propose a novel loss combination variant that improves significantly EDT and C50.
  • We provide convenient and easy access to audio-visual novel view synthesis by releasing a Nerfstudio-compatible module.
@@
Figure 1: NeRAF learns radiance and acoustic fields from a set of images and audio recordings. It synthesizes binaural RIRs and RGB images at novel camera, microphone and source positions and orientations. NeRAF benefits from cross-modal training without requiring co-located audio-visual data and can render spatially separate modalities. NeRAF enables auralization and audio spatialization, along with enhanced image rendering, which are essential for creating a realistic perception of space.
@@ 2 Related works
NeRF [34] synthesizes novel photorealistic views of a scene by modeling it as a continuous function that maps 3D spatial coordinates and viewing directions to radiance. Although it has demonstrated impressive results on objects and small bounded regions, it struggles with complex scenes where the camera can be oriented in any directions and content may exist at any distance. Building upon NeRF, [33] handles unconstrained photo collections, enabling the use of diverse image sets. Other works improved NeRF to learn more complex scenes [2, 62] with higher efficiency and quality [35, 58]. Recently, Nerfstudio [53] provided a modular framework that allows for a simplified end-to-end process of creating, training, and testing NeRF. It combines many existing NeRF improvements to create Nerfacto, a model suited for real-data captures of static scenes. Our method is designed to be a Nerfstudio module bringing novel audio generation to existing state-of-the-art NeRF methods. Moreover, our method improves NeRF performance on large complex scenes given a small image set by leveraging the simultaneous neural acoustic fields training.
@@
Several works have extended neural fields applicability to the audio domain. Luo et al. [30] (NAF) and Su et al. [51] (INRAS) model acoustic propagation in a scene by learning an implicit representation that maps emitter-listener location pairs to binaural RIRs. While NAF conditions the acoustic field by learning local geometric information, INRAS performs audio scene feature decomposition based on the acoustic radiance transfer model. AV-NeRF [28] provides vision-based information to the acoustic field. They rely on pre-trained Nerfacto to render RGB-D images at the microphone pose and extract their features to condition binaural rendering at the same pose. Similarly, our method performs binaural RIR prediction by continuously mapping emitter-listener pairs to the acoustic field. Unlike AV-NeRF, we can decouple the camera and microphone poses used for training. This allows to benefit from unequal amounts of data sources, which is beneficial outside of a simulator. Additionally, at inference, each modality can be rendered independently for greater versatility. We go beyond camera field-of-view information by using NeRF to obtain a 3D grid of the scene filled with color and density information, bringing richer insights to the acoustic field. Furthermore, NeRAF’s cross-modal training benefits both modalities, improving NeRF performances on complex scenes.
@@
RIR have various applications, such as enabling immersive and spatialized audio for AR and VR, performing de-reverberation or providing insights about room acoustics. However, RIR collection in real-world environment is time consuming and needs specialized hardware. Consequently, synthesizing them has been a longstanding research topic. Traditional methods rely on simulated approaches such as wave-based [24, 55, 41, 3] or geometric methods [44, 46, 27, 56, 5, 49]. While these methods effectively simulate sound propagation in space, wave based methods face computation complexity and geometric methods struggle to simulate low-frequency acoustic phenomena such as interference and diffraction. Recent works have proposed learning based methods leveraging the modeling ability of neural networks to learn room acoustics and estimate RIR [31, 30, 51, 42, 43].
@@
Some works rely on both audio and visual information to perform acoustic related tasks such as audio binauralization [20, 22], auralization [8, 48], de-reverberation [10, 14] or RIR prediction [47, 31]. Audio-visual learning has also demonstrated promising results for other tasks such as improving depth prediction [15, 4, 36, 63], performing floorplan reconstruction [32, 40], navigating through a space [61, 21, 13, 7, 19] and estimating poses [13].
Figure 2: Pipeline Overview. NeRF maps 3D coordinates and orientations to density and color. The grid sampler fills a 3D grid representing the scene by querying the radiance field with voxel center coordinates and multiple viewing directions. NAcF learns to map source-microphone poses and directions to STFT coefficients. It is conditioned by extracted scene features. Predicted RIRs can be convolved with audio to obtain auralized and spatialized sound matching the scene.
3 Task
In this study, we focus on simulating acoustic phenomena within a static complex environment based on multiple camera images and audio recordings from known microphone and source poses and orientations. Our goal is to simulate the audio-visual experience for novel microphone-source and camera positions. At inference, our model enables the auditory and visual exploration of the environment from any sensor position without requiring new observations. Consequently, our approach facilitates the synthesis of realistic video paired with matching spatial audio.
@@ 4 Method
Our method is designed as a Nerfstudio module that renders RGB-D images and spatialized audio at any camera and microphone-source pose and orientation. During training, it concurrently learns neural radiance and acoustic fields. At inference, each part can be used independently.
NeRAF pipeline is shown in Figure 2. It comprises 3 modules: NeRF, grid sampler and NAcF. NeRF generates RGB-D images. The grid sampler fills a 3D grid representing the scene with color and density by querying NeRF’s radiance field with the center coordinates of each voxel and multiple viewing directions. NAcF renders binaural audio conditioned by implicit materials and geometric insights provided by the encoded 3D grid.
4.1 Learning scene acoustics using RIRs
To learn the acoustic field of an environment, our method models room impulse responses. RIRs depend on scene geometry and materials. For instance, early reflections are related to the distance from nearby obstacles while late reverberation corresponds to the size and structure of the scene. Learning scene acoustic via RIRs offers several advantages: (1) RIRs capture the acoustic characteristics of a room between emitter-listener positions; (2) by convolving given sounds with RIRs, we can auralize them, effectively synthesizing sounds that mimic the perception of being in the scene; (3) binaural RIRs contain the head-related function transfer (HRTF) necessary to spatialize sound. Thus, by learning to model RIRs, our method captures the sound propagation phenomena necessary for audio synthesis with rich acoustic properties.
@@ 4.2 Neural Radiance Field
Neural radiance field (NeRF) [34] renders photo-realistic images from new view points. It represents a static scene as a 5D vector-valued continuous function whose input is a 3D location $\textbf{X}=(x,y,z)$ and 2D viewing direction $\textbf{d}=(\theta,\phi)$, and output is color $\textbf{c}=(r,g,b)$ and volume density $\sigma$:
@@
NeRF shoots rays $\textbf{r}(t)=\textbf{o}+t\textbf{d}$ through image pixels from camera origin o between the near $t_{near}$ and far $t_{far}$ boundaries of the scene. 3D points $\textbf{X}_{n}=\textbf{o}+t_{n}\textbf{d}$ are sampled along the ray and, along with d used as input of two MLPs. The first one maps $\textbf{X}_{n}$ to the view-independent density $\sigma_{n}$ and a corresponding feature vector. The second MLP takes the feature vector and d to produce view-dependent color $\textbf{c}_{n}$. The final color $C(\textbf{r})$ and depth $D(\textbf{r})$ of an image pixel are rendered using volume rendering equations:

### v2

# NeRAF: 3D Scene Infused Neural Radiance and Acoustic Fields

@@ Abstract
Sound plays a major role in human perception. Along with vision, it provides essential information for understanding our surroundings. Despite advances in neural implicit representations, learning acoustics that align with visual scenes remains a challenge. We propose NeRAF, a method that jointly learns acoustic and radiance fields. NeRAF synthesizes both novel views and spatialized room impulse responses (RIR) at new positions by conditioning the acoustic field on 3D scene geometric and appearance priors from the radiance field. The generated RIR can be applied to auralize any audio signal. Each modality can be rendered independently and at spatially distinct positions, offering greater versatility. We demonstrate that NeRAF generates high-quality audio on SoundSpaces and RAF datasets, achieving significant performance improvements over prior methods while being more data-efficient. Additionally, NeRAF enhances novel view synthesis of complex scenes trained with sparse data through cross-modal learning. NeRAF is designed as a Nerfstudio module, providing convenient access to realistic audio-visual generation. Project page: https://amandinebtto.github.io/NeRAF
Preprint. Under review.
@@ 1 Introduction
Sound is fundamental to human perception. Imagine standing in the middle of a bustling city. It’s not only the visual cues but the sounds of car horns, footsteps, and chatter that guide your perception and decisions. Beyond conveying environmental details, sound provides crucial insights into context and atmosphere, enriching our understanding of the world in ways that sight alone cannot. In gaming and AR/VR simulations, audio is key to immersion, enhancing the realism of virtual environments and making experiences feel authentic.
@@
Advances in novel view synthesis have enabled the generation of high-quality, photorealistic images from any camera position using a set of captured images (Mildenhall et al., 2020; Martin-Brualla et al., 2021; Zhang et al., 2020; Tancik et al., 2023), unlocking exciting applications for simulated environments. Yet, these methods lack acoustic synthesis, which is essential for fully immersive experiences. Learning scene acoustics presents additional challenges (Luo et al., 2022). As sound travels from its source to the listener it is influenced by the geometry of the space and the materials of objects and surfaces. Room Impulse Responses (RIRs) capture these complex interactions for specific source-listener pairs, but collecting RIRs is a laborious process, requiring sounds to be played and recorded at multiple positions.
To address this, recent research has focused on estimating RIRs at novel positions from sparse data (Majumder et al., 2022; Singh et al., 2021; Somayazulu et al., 2024). Inspired by NeRF’s success in image synthesis, emerging studies (Luo et al., 2022; Su et al., 2022; Liang et al., 2023a; b) apply neural fields to learn the acoustic properties of an environment, allowing for audio synthesis from new source and microphone positions. However, these approaches often overlook the critical influence of 3D scene geometry on acoustics (Luo et al., 2022; Su et al., 2022; Liang et al., 2023a), or they rely on additional annotations that are unavailable in real-world scenarios (Liang et al., 2023b).
In response, we introduce NeRAF, a method that generates both novel views and RIRs at new sensor positions by jointly learning radiance and acoustic fields (Figure 1). NeRAF queries the radiance field to construct a voxel representation of the scene that encodes radiance and density. This grid conditions the acoustic field with appearance and geometric 3D priors, without the need for additional annotations. NeRAF’s cross-modal approach benefits both modalities: it achieves state-of-the-art RIR synthesis while being more data-efficient and enhances novel view generation in complex scenes. We validate our method on Sound Spaces (Chen et al., 2020a; 2022b), a simulated dataset, and on RAF (Chen et al., 2024), a real-world dataset. Designed as a Nerfstudio module, NeRAF offers convenient access to audio-visual generation necessary for immersive experiences. Although the approach is cross-modal, each modality can be generated independently and the model can be trained without co-located audio and visual sensors, providing flexibility for various applications. We release NeRAF’s code to the community.
@@
Figure 1: NeRAF synthesizes audio-visual data at novel sensor positions by learning radiance and acoustic fields from a collection of images and audio recordings. It enables audio auralization and spatialization, as well as improved image rendering, all of which are crucial for creating a realistic perception of space. NeRAF leverages cross-modal learning without the need for co-located audio and visual sensors for training. Our method allows for the independent rendering of each modality.
@@ 2 Related works
NeRF (Mildenhall et al., 2020) synthesizes novel photorealistic views of a scene by modeling it as a continuous function that maps 3D spatial coordinates and viewing directions to radiance. Although it has demonstrated impressive results on objects and small bounded regions, it struggles with complex scenes where the camera can be oriented in any directions and content may exist at any distance. Building upon NeRF, (Martin-Brualla et al., 2021) handles unconstrained photo collections, enabling the use of diverse image sets. Other works improved NeRF to learn more complex scenes (Barron et al., 2022; Zhang et al., 2020) with higher efficiency and quality (Müller et al., 2022; Xu et al., 2022). Recently, Nerfstudio (Tancik et al., 2023) provided a modular framework that allows for a simplified end-to-end process of creating, training, and testing NeRF. It combines many existing NeRF improvements to create Nerfacto, a model suited for real static scenes. NeRAF is designed as a Nerfstudio module, bringing novel audio synthesis to existing state-of-the-art NeRF methods. Its cross-modal learning improves Nefacto performance on large complex scenes given a small image set.
@@
Several works have extended the applicability of neural fields to the audio domain. Luo et al. (2022) (NAF) and Su et al. (2022) (INRAS) model acoustic propagation in a scene by learning an implicit representation that maps emitter-listener positions to RIRs. NAF conditions the acoustic field by learning local geometric information, while INRAS performs audio scene feature decomposition based on the acoustic radiance transfer model. NACF (Liang et al., 2023b) and AV-NeRF (Liang et al., 2023a) provide vision-based information to the acoustic field. NACF assumes the pre-acquisition of RGB-D images along with acoustic coefficients of surfaces. However, the latter is unavailable in real-life scenarios. AV-NeRF relies on a pre-trained Nerfacto model to render RGB-D information at the microphone’s position. It then conditions the audio rendering at the same pose. In contrast, NeRAF extends beyond the camera’s field of view by using NeRF to capture 3D scene information – without the need for more annotations – providing richer insights for the acoustic field. Contrary to AV-NeRF, our method does not need co-located audio and visual sensors for training. Therefore, it handles unequal amounts of training data for each modality, which is beneficial in real-world settings. Additionally, at inference, the vision modality is not necessary for audio rendering leading to greater versatility.
@@
RIR have various applications, such as enabling immersive and spatialized audio for AR and VR, performing de-reverberation or providing insights about room acoustics. However, RIR collection in real-world environment is time consuming and needs specialized hardware. Consequently, synthesizing them has been a longstanding research topic. Traditional methods rely on simulated approaches such as wave-based (Gumerov & Duraiswami, 2009; Thompson, 2006; Raghuvanshi et al., 2009; Bilbao et al., 2019) or geometric methods (Savioja & Svensson, 2015; Schissler & Manocha, 2016; Krokstad et al., 1968; Vorländer, 1989; Cao et al., 2016; Sprunck et al., 2022). While these methods effectively simulate sound propagation, wave-based methods face computation complexity and geometric methods struggle to simulate low-frequency acoustic phenomena such as interference and diffraction. Recent works have proposed to leverage the modeling ability of neural networks to learn room acoustics and estimate RIRs (Majumder et al., 2022; Luo et al., 2022; Su et al., 2022; Ratnarajah et al., 2021; 2022).
@@
Some works rely on both audio and visual information to perform acoustic related tasks such as audio binauralization (Gao & Grauman, 2019; Garg et al., 2023), auralization (Chen et al., 2022a; Somayazulu et al., 2024), de-reverberation (Chen et al., 2023a; Chowdhury et al., 2023) or RIR prediction (Singh et al., 2021; Majumder et al., 2022). Audio-visual learning has also demonstrated promising results for other tasks such as improving depth prediction (Christensen et al., 2020; Brunetto et al., 2023; Parida et al., 2021; Zhu et al., 2024), performing floorplan reconstruction (Majumder et al., 2023; Purushwalkam et al., 2021), navigating through a space (Younes et al., 2023; Gao et al., 2020; Chen et al., 2023c; 2020b; Gan et al., 2020) and estimating poses (Chen et al., 2023c).
Figure 2: NeRAF overview. NeRF maps 3D coordinates, $\mathbf{X}$, and orientations, $\mathbf{d}$, to density and color. The grid sampler fills a 3D grid representing the scene by querying the radiance field with voxel center coordinates and multiple viewing directions. NAcF learns to map source-microphone poses and directions to STFT. It is conditioned by extracted scene features. Predicted RIRs can be convolved with audio to obtain auralized and spatialized audio matching the scene.
@@ 3 Method
NeRAF generates both RGB images and binaural RIRs for any camera and microphone-source position and orientation. It achieves this by jointly learning neural radiance and acoustic fields. Although trained together, the audio and visual components can be used independently.
The NeRAF pipeline, shown in Figure 2, consists of three modules: NeRF, the grid sampler, and NAcF. The NeRF learns the radiance field and generates RGB images. The grid sampler queries NeRF’s radiance field with each voxel’s center, filling the 3D grid with color and density. NAcF then renders binaural RIRs, conditioned by the implicit appearance and geometric 3D priors derived from the features extracted from this 3D grid.
3.1 Modeling Acoustic through RIRs
To model the acoustic of an environment, our method learns RIRs. They describe how sound propagates between one emitter and listener depending of the scene’s geometry and material properties. For example, early reflections provide information about the distance to nearby surfaces, while late reverberation reflects the overall size and structure of the scene. Learning acoustics through RIRs offers several key benefits: (1) RIRs encapsulate the acoustic behavior between specific emitter and listener positions; (2) By convolving audio signals with RIRs, we can synthesize sound that mimics how it would be heard within the scene; (3) Binaural RIRs include head-related transfer functions (HRTFs), which are essential for accurately spatializing audio.
By learning to generate RIRs, our method captures essential sound propagation phenomena, enabling realistic audio synthesis with rich acoustic details.
@@ 3.2 Neural Radiance Field
Neural radiance field (NeRF) (Mildenhall et al., 2020) renders photo-realistic images from new view points. It represents a static scene as a 5D vector-valued continuous function whose input is a 3D location $\textbf{X}=(x,y,z)$ and 2D viewing direction $\textbf{d}=(\theta,\phi)$, and output is color $\textbf{c}=(r,g,b)$ and volume density $\sigma$:
@@
NeRF shoots rays $\textbf{r}(t)=\textbf{o}+t\textbf{d}$ through image pixels from camera origin o between the near $t_{near}$ and far $t_{far}$ boundaries of the scene. 3D points $\textbf{X}_{n}=\textbf{o}+t_{n}\textbf{d}$ are sampled along the ray and, along with d used as input of two MLPs. The first one maps $\textbf{X}_{n}$ to the view-independent density $\sigma_{n}$ and a corresponding feature vector. The second MLP takes the feature vector and d to produce view-dependent color $\textbf{c}_{n}$. The final color $C(\textbf{r})$ of an image pixel are rendered using volume rendering equations:
@@
NeRAF builds on Nerfacto (Tancik et al., 2023) as its NeRF backbone. Nerfacto integrates advancements from several works following (Mildenhall et al., 2020) including camera pose refinement (Wang et al., 2021; Lin et al., 2021), per image appearance conditioning (Martin-Brualla et al., 2021), proposal sampling (Müller et al., 2022), scene contraction and hash encoding. We selected Nerfacto for its robust performance in handling real-world static scenes. However, our method is NeRF-agnostic as it relies directly on the radiance field without altering the underlying NeRF model.
3.3 Grid Sampler
Figure 3: Grid sampler. We represent the scene as a grid of voxels. The grid is populated by querying the radiance field with the center coordinates of each voxel, $\textbf{X}_{vi}$, along with multiple viewing directions, $\textbf{d}_{1\rightarrow N}$. We average the color values obtained from these directions. It results in a 7-channels 3D grid containing color $\mathbf{\widehat{C}}_{vi}$, density $\sigma_{vi}$ and the 3D coordinates of the voxel centers.
Sound propagation is omnidirectional and influenced by the 3D geometry of the environment. Therefore, we provide our neural acoustic field with 3D geometric and appearance priors about the scene.
To this end, we use our grid sampler, illustrated in Figure 3, to construct a 3D volume from the NeRF model. The scene is represented as a 3D grid of $S^{3}$ voxels with coordinates between 0 and 1 in accordance with NeRF’s scene contraction. We shows in Section 4.3, that $S=128$ leads to good performances and use it for all our experiments.
We populate the grid by sequentially selecting batches of voxel centers $\textbf{X}_{vi}$ as queries to the radiance field. By processing voxels in order, we ensure that the grid is progressively updated to reflect the improvements in the NeRF model. To avoid slowing down training, we do not update the entire grid at each training iteration and process in batches. This sampling method guarantees complete grid population.
As NeRF color is view-dependent, we follow (Chen & Lee, 2023) and form $N=18$ viewing rays per $\textbf{X}_{vi}$. The goal is to obtain an average color of the voxel. The radiance field returns a density value $\sigma_{vi}$ and a color value $\textbf{c}_{vi_{j}}$ per $j\in[1,N]$ viewing directions:
$\text{NeRF}:(\textbf{X}_{vi},\textbf{d}_{1\rightarrow N})\rightarrow(\textbf{c}_{vi_{1\rightarrow N},\sigma_{vi}}).$ (4)
For color, we apply the volume rendering equation to a ray containing a single point, $\textbf{X}_{vi}$, with each viewing direction and average the values across these directions to obtain Ĉ$_{vi}$. For density, we compute the alpha compositing value $\alpha=1-\exp(-\sigma_{vi}\delta)$ where $\delta$ is a chosen small value. We populate the grid with these values and the 3D coordinates of the voxels, resulting in a 7-channel representation.
3.4 Neural Acoustic Field
The goal of our neural acoustic field (NAcF), presented in Figure 4, is to learn a continuous neural representation of the scene’s acoustics. To this end, NAcF maps microphone coordinates $\textbf{X}_{m}=(x_{m},y_{m},z_{m})$ and orientations $\textbf{d}_{m}=(\theta_{m},\phi_{m})$ along with source position $\textbf{X}_{s}=(x_{s},y_{s},z_{s})$ and orientation $\textbf{d}_{s}=(\theta_{s},\phi_{s})$ to the short-time Fourier transform (STFT) of a room impulse response:
$\text{NAcF}:(\textbf{X}_{m},\textbf{d}_{m},\textbf{X}_{s},\textbf{d}_{s},t)\rightarrow\text{{RIR}}(f,t).$ (5)
We learn RIR through their STFT representation, resulting in $\text{STFT}\in\mathbb{R}^{C\times F\times T}$, where C, F and T denote the number of channels, frequency bins, and time bins, respectively. This is motivated by the fact that the time-frequency representation is smoother and more suitable to neural network prediction than the time domain (Luo et al., 2022). Thus, NAcF learns to infer log-magnitude STFTs. To better capture perceptual effects such as reverberation, we query NAcF with STFT time bins $t$. Consequently, NAcF outputs the $F$ frequency bins corresponding to the time query. We normalize $t$ between $[0,1]$ and encode it with positional encoding. Similar to NeRF (Mildenhall et al., 2020) pose, microphone and source coordinates as well as direction are also embedded to a higher dimensional space using respectively positional encoding and spherical harmonic encoding.
Given the voxel grid, we concurrently train a ResNet3D (He et al., 2016) to extract relevant features for sound propagation. They serve as implicit indicators of scene’s 3D geometry, which is central for acoustic modeling. Moreover, as shown in previous works (Chen et al., 2023b; Singh et al., 2021), information about materials can be inferred by a ResNet architecture from the scene’s appearance and contribute to audio synthesis. We perform multimodal fusion by concatenating the features vector from the grid to the encoded poses, directions and the time query. Similar to NeRF, NAcF comprises two MLP blocks. The first block maps the input to a feature vector that embeds the input information and the acoustic scene. The second focuses on learning microphone specifics. For binaural microphones, this block learns the HRTF necessary for audio spatialization. It contains one MLP per microphone channel.
To render the complete STFT, we query all time bins $t\in[0,T]$. Finally, we use Griffin-Lim (Griffin & Lim, 1984; Perraudin et al., 2013) algorithm to obtain RIR waveforms from magnitude STFTs.
Figure 4: Neural acoustic field. NAcF maps microphone-source poses and directions to either binaural or monaural RIRs, with the number of output heads adjusted accordingly. NAcF is conditioned on scene features extracted through a 3D ResNet. Both poses and time queries are transformed into a higher-dimensional space using positional encoding, while directions use spherical harmonic encoding. The output of NAcF is a vector containing $F$ frequency values for each time query, $t$.
3.5 Learning Objective
Cross-modal learning has proven beneficial in multiple works (Section 2). Thus, NeRAF trains jointly NeRF and NAcF. NAcF training is slightly delayed as it begins when the grid has been updated a few times. The feature extractor is jointly trained with the radiance and acoustic fields.
We do not modify the NeRF loss function $\mathcal{L}_{V}$. It usually consists in an MSE loss between ground-truth pixel color $C(\textbf{r})$ and the rendered one $\hat{C}(\textbf{r})$, and complementary losses, $\mathcal{L}_{\text{comp}}$, specific to each NeRF.